OpenStudy (anonymous):

In the gym, there is a scale that measures people's weight. The needle is 22 cm long and rotated 124 degrees when Evelyn stepped on it. How far did the needle travel as it rotated?

OpenStudy (anonymous):

1364π 2728π 682pi/45 45pi/682

OpenStudy (anonymous):

@campbell_st @ranga

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@amistre64

hartnn (hartnn):

so basically you need length of arc \(s= r\theta \) first convert 124 degrees to radians because \(\theta \) there needs to be in radians \(124^o = 124\times \dfrac{\pi^r}{180^o} = ... \)

hartnn (hartnn):

then you just multiply that answer by 22

OpenStudy (anonymous):

so 110.88?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

please keep \(\pi \) as \(\pi\) only just simplify 22*124/180

OpenStudy (anonymous):

so 15.15?

OpenStudy (anonymous):

what do i do next? @hartnn

hartnn (hartnn):

which is also 682/45

OpenStudy (anonymous):

OMG! thank you!!! so I don't have to use the pi?

OpenStudy (cwrw238):

1 complete circle = 2 pi r radians = 360 degrees for 124 degrees distance travelled = (124 / 360) * 2 pi r plug in r = 22 which simplifies to 682pi/45

hartnn (hartnn):

we kept pi as pi \(22\times 124\pi/180 = 681\pi /45\)

hartnn (hartnn):

682***

OpenStudy (anonymous):

Okay that totally makes more sense thank you!

hartnn (hartnn):

welcome ^_^